OpenStudy (anonymous):

@iambatman @KingGeorge @hartnn @jim_thompson5910 @beccaboo333 @jdoe0001 @mathmale @Awesome781 @geerky42

OpenStudy (anonymous):

Can you explain this really simple? I need it step by step and don't move onto the next step unless I understand it. I would prefer if you could use the drawing tool too. :) Please and Thank You! Also please don't give the answer I wanna learn how to work it out. Find the solution of the linear equation. 2/3 (6 - 3x) = (1 - x) + (3 - x) A. x = -8 B. x = 8 C. infinitely many solutions D. no solution

geerky42 (geerky42):

Let start with simplifying LHS (left hand side) Can you do it?

OpenStudy (anonymous):

Can you show me how using the drawing tool? Please?

geerky42 (geerky42):

Use distributive property: